OpenStudy (anonymous):

a rectangle has an area of 60 square inches and a perimeter of 38 inches. what is the length and width of the rectangle?

OpenStudy (anonymous):

l*w = 60 2l + 2w = 38 solve for l and w

OpenStudy (anonymous):

l*w=60----1 2(l+w)=38 l+w=38/2 l+w=19 l=19-w-----2 substitute 2 in 1 (19-w)(w)=60 19w-w^2=60 w^2-19w+60=0 w^2-15w-4w+60=0 w(w-15)-4(w-15)=0 (w-15)(w-4)=0 w=15,w=4 so length=15 and width=4